OpenStudy (anonymous):

Solve system using substitution 3x+y=5 2x-5y=9

OpenStudy (anonymous):

does anybody know the answer

OpenStudy (anonymous):

( 2, -1 )

OpenStudy (anonymous):

can you show me how you got that plz

OpenStudy (anonymous):

you first get one of the variable alone. I got the y in the 1st eq alone: y=5-3x Then you take the y out of the second eq and put the 5-3x in. Like : 2x-5(5-3x)=9 So u slove it X=2 Then plug x in one of the eqs. For example : 3*2+y=5 6+y=5 y=-1 So the solution is ( 2 , -1 )

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcome !!

OpenStudy (anonymous):

can you help me on the other math problems i have. i don't want to be a bother

OpenStudy (anonymous):

yes. what is it ? :)

OpenStudy (anonymous):

y-2=x -x=y

OpenStudy (anonymous):

For that one i am not sure how to do it... sorry :P coz when i plug it in, the solution didnt work for one of the eq ...

OpenStudy (anonymous):

Try ask somebody else

OpenStudy (anonymous):

its ok how about 6y+4x=12 -6x+y=-8

OpenStudy (anonymous):

i got (1.5,2) but when i check it i got it wrong

OpenStudy (anonymous):

its (1.5, 1 )

OpenStudy (anonymous):

oh i know what i did i misplaced a number thanks

OpenStudy (anonymous):

no probbb :D

OpenStudy (anonymous):

how about 2y-3x=4 x=-2

OpenStudy (anonymous):

(-2 , -1 )

OpenStudy (anonymous):

can you show your work plz

OpenStudy (anonymous):